A valószínűségszámítás I. gyakorlaton való részvétel

dátum:  I. 29. II. 5 . II. 12. II. 19. II. 26. III. 5. III. 12. III. 19

diákok:

Barabás Timea  +  +  −  +  +  −  +  +
Bíró Ildikó  +  −  +  −  +  +  −  +
Bulátki János  +  −  +  −  +  +  +  −
Csaba Anikó  +  −  +  +  +  +  +  +
Dorogi Edina  +  +  −  −  +  +  +  +
Jámbor Tamás  +  +  +  +  +  +  +  +
Hubert Balázs  −  −  +  +  +  +  −  −
Kaponyás Ágnes  +  −  +  −  +  +  −  +
Lefter Simon  −  +  −  −  −  −  −  −
Medve Diana  −  −  +  +  +  −  −  +
Mészáros Fruzsina  +  +  +  +  +  +  +  +
Nagy Adrienn  +  −  +  −  −  +  −  +
Öldi Kinga  −  +  +  +  +  +  +  +
Pálfi Anita Katalin  +  −  −  −  −  −  −  −
Suta Judit  +  +  +  +  +  −  −  −
Szalamics Zoltán  +  +  +  −  +  +  +  −
Szűcs Orsolya  +  −  −  +  +  −  +  +
Tanka Irén  +  −  +  +  +  +  +  +
Tóth Boglárka  +  +  +  +  +  +  −  +
Tóth György  +  +  +  +  +  +  −  +

Furó Attila: III. 12

dátum:  IV. 9. IV. 16 . IV. 23. IV. 30. V. 7.

diákok:

Barabás Timea  +  +
Bíró Ildikó  −  −
Bulátki János  +  −
Csaba Anikó  +  +
Dorogi Edina  −  +
Jámbor Tamás  +  +
Hubert Balázs  −  +
Kaponyás Ágnes  −  −
Medve Diana  −  −
Mészáros Fruzsina  +  +
Nagy Adrienn  +  −
Öldi Kinga  +  +
Suta Judit  −  +
Szalamics Zoltán  +  −
Szűcs Orsolya  +  −
Tanka Irén  −  −
Tóth Boglárka  +  −
Tóth György  −  −